OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3). y = 2x − 7 y = -1/2x − 2 y = 2x − 2 y = −1/2x − 7

OpenStudy (anonymous):

In order for a line to be parallel to y=2x+2, it must have the same slope. The coefficient (number) in front of x represents the slope of the line, in this case, 2. Thus the answer can be narrowed down to y=2x-7 or y=2x-2. To find which of these lines passes through (5,3), plug the x coordinate into each equation like so: y=2(5)-7 y=2(5)-2 y=10-7 y=10-2 y=3 y=8 since the y coordinate in (5,3) is 3, the equation y=2x-7 passes through the point (5,3)

OpenStudy (anonymous):

I'm not sure whats the answer

OpenStudy (anonymous):

wait, so it's a?

OpenStudy (anonymous):

y=2x-7 would be the answer because it is has the same slope as the given line and passes through (5,3)

OpenStudy (anonymous):

alright thanks bro